OpenStudy (yonatanbesrat):

x^2-6x+9=0 a,1 real solution x^2-3x+15=0 b,2 real solution c,no real solution x^2-8x+7=0

OpenStudy (3mar):

If I may help?

OpenStudy (3mar):

@Yonatanbesrat

OpenStudy (metalslayer):

For the last number, (9, 15 and 7) is there a variable next to them? If they all have a variable here is the answer: So the answer will be zero for all variables, because all of the numbers have a coefficient and they are all equal to zero.

OpenStudy (metalslayer):

That's one answer.

OpenStudy (holsteremission):

Check the discriminant: \[ax^2+bx+c=0\implies \Delta=b^2-4ac\]There are three cases to consider: \[\begin{cases}\Delta>0\implies 2\text{ real solutions}\\[1ex]\Delta=0\implies1\text{ real solution}\\[1ex]\Delta<0\implies 2\text{ complex solutions}/0\text{ real solutions}\end{cases}\]